OpenStudy (anonymous):

Find the Vertex, Value of P, Axis of Symmetry, Focus, Directix, then graph the equation: y + 2 = 1/4(x-1)^2

OpenStudy (anonymous):

If I could also get help on this question, it would be much loved. These are review questions for an upcoming exam and I would appreciate all the help I could get on it: The shape of a park can be modeled by a circle with the equation x^2 + y^2 = 1600. A stretch of highway near the park is modeled by the equation y = 1/40(x - 40)^2. At what points does a car on the highway enter or exit the park?

OpenStudy (campbell_st):

ok... so the standard form of the parabola I use is (x - h)^2 = 4a(y - k) the vertex is (h, k) and a = focal length... so you're equation can be written as \[4(y + 2) = (x - 1)^2\] can you identify the vertex and focal length a ?

OpenStudy (anonymous):

@campbell_st - I figured it out, thanks! But I still need help with the other question I mentioned. Do you think you could help me out with that one? And also this one if you can: 4cosθ = 1, for 0° ≥ θ ≥ 360°

OpenStudy (anonymous):

I had figured it out ahead of time on my own. But I'm still stuck on the other two.

OpenStudy (campbell_st):

ok... so divide both sides by 4 \[\cos(\theta) = \frac{1}{4}\] given the domain, there are 2 answers... 1st quadrant and 4th quadrant.... this is where cos is positive so find the 1st quadrant angle... and 4th quadrant is \[360 - \theta\] hope it helps

OpenStudy (anonymous):

@campbell_st can you tell me how to figure out the quadrants? θ = 1/4, so would we do cos(1/4) to figure out the answer?

OpenStudy (anonymous):

I entered it in my calculator and got this: θ = 2πn - cos^-1(1/4) Is it right?

OpenStudy (campbell_st):

well there is this diagram  it comes from the unit circle... the diagram shows which trig ratios are positive in each quadrant...

OpenStudy (campbell_st):

so is cos(@) = 1/4 whats the value of @ normally its just shift cos (1/4) =

OpenStudy (anonymous):

So it's 360° - 1/4?

OpenStudy (anonymous):

@campbell_st I got this answer after putting it in my calculator but I don't know if it's right: θ = 2πn - cos^-1(1/4)

OpenStudy (campbell_st):

no if \[\cos(\theta) = \frac{1}{4} ....then..... \theta = \cos^{-1}(\frac{1}{4}) = 75.5^o\] does that make sense... and that's the 1st quadrant angle

OpenStudy (anonymous):

@campbell_st - So then would the fourth quadrant be 224.5°?

OpenStudy (campbell_st):

no... its 360 - 75.5

OpenStudy (anonymous):

@campbelll_st - I mean 284.5°

OpenStudy (campbell_st):

that makes more sense